Compute the statistical power Type II error rate 0.25 sample size 100 alpha level 0.01
Answer #1

Given that,

Type II error rate=0.25

Sample size=100

Significance level=0.01=α

Now, Type II error = β = 0.25 (Given)

Power = 1 - Type II error =1 - 0.25 = 0.75 (Ans)